# Observing network dynamics through sentinel nodes

### v1

@@ II.4 Weighted averaging
We find that the additional optimization of node weights leads to substantially smaller approximation error, $\varepsilon$, than the unweighted average examined in the previous sections. Across all dynamics and networks, $\varepsilon$ is on average an additional 46 times smaller when we optimize both node set and node weight than when we only optimize the node set (see SI section S8 for the statistical results).
However, there is no extra advantage to optimizing node weights when one evaluates $\varepsilon$ on alternative dynamics. As in section II.3, we generated, for each dynamics and network, 100 node sets with their node selection and weights optimized on a training dynamics. We then evaluated $\varepsilon$ of the optimized node set on each other test dynamics. We found that, when evaluated on the test dynamics, $\varepsilon$ was not significantly different between the optimized node sets with additional weight optimization and those without ($p=0.591$). Therefore, although optimizing node weights drastically improves the approximation to $\overline{x}$ if we know the actual dynamics, this advantage is lost if we do not know it.
@@ III Discussion
As our understanding of complex system dynamics advances [3, 1, 25, 26, 27, 28, 29], a crucial bottleneck that we continue to face is our limited empirical access to the states of their multitude of relevant parameters. This bottleneck hinders the development of new theoretical analyses and empirical validation of existing theories on dynamics. A similar challenge in biology and medicine has been that it is often difficult to track every single metabolite, protein, or cell. In that context, the identification of biomarkers [30, 31] can offer dramatic breakthroughs, potentially allowing us to identify the state of the entire system just by tracking a small set of indicative parameters. Our work here envisions an equivalent opportunity in the context of complex networks.
@@
A crucial aspect of our formulation is that the detected sentinels are largely agnostic to the dynamic model. This indicates that one can transfer the learned set of nodes from one dynamics to infer the state of the system under another dynamics. Such transfer learning is especially useful for systems with unknown dynamics, capturing a rather common challenge in the context of complex networked systems [32, 33, 34, 35, 36]. To further capitalize on this potential advantage, future work should aim to extract a master model, i.e., a nonlinear system of equations whose sentinel nodes optimally transfer to other types of dynamics.
@@
Our work differs from other approaches for approximating the population activity of or efficiently observing a network by attempting to accurately estimate $\overline{x}$. First, there are methods to select a small number of sentinel nodes for constructing early warning signals for anticipating regime shifts [16, 17, 18, 19, 20], but there is no a priori reason to consider that these sentinel nodes are also good at representing $\overline{x}$. Second, the theory of network observability allows us to determine the node set, observation of which enables one to reconstruct the system’s complete activity [37]. The method has also been applied for selecting sentinel nodes with which to construct early warning signals [18]. However, this observability concerns inference of the initial condition, i.e., the activity of each node at time 00. In fact, $\overline{x}$ in the equilibrium or the time course of it, rather than its initial value, would be of practical interest in applications.
Our work also differs from prior work in the importance it attributes to hub nodes in heterogeneous networks. Both the GBB [9] and DART [10, 11] reduce dynamical systems on networks into a similar dynamical system of small dimensions. A key idea behind these methods is to theoretically find an appropriate linear combination $\langle x\rangle=\sum_{i=1}^{N}a_{i}x_{i}^{*}$ of $x_{1}^{*}$, $\ldots$, $x_{N}^{*}$ and the dynamical system that $\langle x\rangle$ approximately obeys. To do so, GBB and DART use the degree vector and the dominant eigenvector of the adjacency matrix as $\bm{a}=(a_{1},\ldots,a_{N})$, respectively, as the first $\langle x\rangle$. In fact, many large empirical networks are scale-free networks, i.e., those in which the degree obeys an approximate power-law distribution [38, 39, 40], so that a small fraction of hub nodes have very large degrees. In a related vein, many empirical and model networks show eigenvector localization phenomena, with which the most entries of the dominant eigenvector of the adjacency matrix are close to 00 [41, 42]. Eigenvector localization implies that, at least in large degree-heterogeneous but otherwise uniformly random networks, most of the entries of $\bm{a}$ are close to 00 and do not contribute to $\langle x\rangle$. Therefore, $\langle x\rangle$ in the GBB and DART is approximately a weighted sum of a small number of $x_{i}$. This situation apparently resembles selecting $n$ ($\ll N$) nodes for $S$ in our framework. However, these methods and our sentinel node approximation select very different nodes. The nodes effectively used by GBB and DART (i.e., those whose $a_{i}$ values are far from 00) tend to be hub nodes. Therefore, GBB and DART deliberately neglect the behavior of a majority of nodes, which have a small degree. In contrast, our optimized node set tends to avoid hub nodes, which are outliers, and instead mixes nodes with different degrees that are not extremely large. This discrepancy is because our goal is to observe the unweighted average, $\overline{x}=\sum_{i=1}^{N}x_{i}^{*}/N$, while the goal of GBB and DART is to derive a dynamical system approximately closed in terms of $\langle x\rangle$. Although hubs play important roles in dynamics on networks such as impacting on the epidemic threshold [43, 44] and having different responsiveness to perturbation compared to low-degree nodes [25], our optimized node sets does not include them to approximate $\overline{x}$ because their behavior is in general different from that of the majority of nodes.
@@
Our framework accommodates directed and weighted networks, noisy dynamics, and, as mentioned above, the case in which dynamical system equations are unknown. Furthermore, without essential changes, our method allows any weighted sum of $\{x_{1}^{*},\ldots,x_{N}^{*}\}$ as the target quantity to be approximated. Together, these features make our method versatile enough for potential applications. For example, in ecosystem monitoring, our algorithm could help focus monitoring efforts on a small number of key species [45], which may not be the most numerous or well-connected species [46, 47]. Assuming that an estimated network of species interactions is available, which is often the case [48], one can derive a small number of sentinel species (i.e., nodes) by simulating a “ground truth” using a mutualistic species dynamics model on the measured network. Then, the populations of the chosen sentinel species are expected to provide a good approximation of population dynamics of the entire ecosystem.
@@
From a methodological perspective, our drastic reduction from $N$ to just a small number of nodes is rooted in a combined toolbox that mixes theoretical modeling with machine learning techniques. The theoretical models, captured within the low-dimensional dynamical systems derived by GBB, DART, or other methods, represent our analytical insights into the system’s internal mechanisms. Such insights, however, are limited by the prohibitive complexity of the system. This is precisely where the machine learning component helps complement the theoretical approach, lending its computational power to help simplify the scale of the problem. This synergy between theoretical insights and the predictive power of modern computational learning technology represents, in our view, one of the main avenues towards predicting, observing, and influencing complex network behavior.
@@ IV.1 Node Selection
We seek a set $S$ of sentinel nodes with $|S|=n$, $n\ll N$. We assume that each node in $S$ has equal weight in determining $\overline{x}^{\prime}$, as in Eq. (2). We relax this assumption in section II.4. We quantify the discrepancy between $\overline{x}$ and $\overline{x}^{\prime}$ by a normalized mean squared error over a given range of the control parameter, i.e., $\varepsilon=\frac{\sum_{\ell=1}^{L}(\overline{x}^{\prime}_{\ell}-\overline{x}_{\ell})^{2}}{L\sum_{\ell=1}^{L}\overline{x}_{\ell}}$, called the approximation error. Subscript $\ell\in\{1,\ldots L\}$ specifies a value of an evenly spaced control parameter, and $x_{i,\ell}^{*}$ represents the equilibrium value of $x_{i}$ for the $\ell$th value of the control parameter; $\ell$ is omitted in the following text when there should be no confusion. We choose a control parameter which, when varied over the first and last (i.e., $L$th) values, causes bifurcations in $x_{i}^{*}$. We set $L=100$ in our numerical experiments.
Note that two other known methods for dynamics reduction on networks, GBB [9, 23, 15] and DART [10, 11, 13, 24, 14], attempt to match a different weighted sum of $x_{1}$, $\ldots$, $x_{N}$ from $\overline{x}$. Specifically, the observables of the GBB and DART are in the form $\sum_{i=1}^{N}a_{i}x_{i}$, where the $a_{i}$ is proportional to the degree of the $i$th node in the case of the GBB and the $i$th entry of the leading eigenvector of the adjacency matrix in the case of the DART.
@@ IV.2 Dynamical system models
where $x_{i}$ represents the probability that the $i$th node is infectious, $\lambda$ is the infection rate, and $\mu$ is the recovery rate [44]. The second term on the right-hand side expresses the rate at which the $j$th node infects the $i$th node. We set $\mu=1$ without loss of generality; changing $(\lambda,\mu)$ to $(c\lambda,c\mu)$ with a constant $c$ is equivalent to scaling the time by a factor of $c$, so it does not affect the equilibrium. We use $\lambda$ as the control parameter in the range $[0,1]$. One obtains $x_{i}=0$ $\forall i$ when $\lambda$ is below a value called the epidemic threshold. Above the epidemic threshold, all $x_{i}^{*}$ ($<1$) values are positive given a positive initial value. The parameter $\lambda$ has the same role as $D$ in the other dynamics models.
@@
To compute equilibrium values of $x_{i}$ at each control parameter value, we initially set each $x_{i}$ to a model-dependent standard value (coupled double-well: $x_{i}=1$, SIS: $x_{i}=0.01$, mutualistic species: $x_{i}=0.001$, gene-regulatory: $x_{i}=2$). Then, we solved the ODEs using the LSODA algorithm [56] provided by the deSolve package for R [57]. We used a total simulation time of $T=15$, which we found was sufficient to allow each of the above dynamics on each network described below to relax to a point where no further change was noticeable. We used the value of each $x_{i}$ at $T=15$ as $x_{i,\ell}^{*}$.
@@ IV.3 Networks
We chose five empirical and five model networks to test our method. All networks were coerced to be undirected, simple (i.e., no self- or multi-edges), and unweighted. We used the largest connected component.
@@
We used eight other networks as well as the dolphin and BA networks in the ANOVA analyses. See SI section S1 for the description of these eight networks.
@@ IV.4 Optimization of node weights
In section II.4, we approximate $\overline{x}_{\ell}$ with a weighted average of the node states in $S$, $\overline{x}^{\prime}_{\ell}=\sum_{i=1;i\in S}^{N}a_{i}x_{i,\ell}^{*}$. The GBB and DART assign weights to all the nodes in the network. In contrast, here we determine weights only for the nodes in $S$ by a quadratic programming optimization.
@@
[…]
@@
Data and code availability
Data and analysis code are available at https://github.com/ngmaclaren/reduction.
Acknowledgements
This work was performed in part at Center for Computational Research, the State University of New York at Buffalo.
@@ Acknowledgements
Funding
B.B. was supported by the Israel Science Foundation (grant no. 499/19), the Israel-China ISF-NSFC joint research program (grant no. 3552/21), and by the VATAT grant for data science research. N. Masuda was supported by the Japan Science and Technology Agency (JST) Moonshot R&D (under grant no. JPMJMS2021), the National Science Foundation (under grant no. 2052720), and JSPS KAKENHI (under grant nos. JP 21H04595, 23H03414, and 24K14840).
@@
Approximating nonlinear dynamics on networks by sentinel nodes
@@
In the main text, we have shown results from two undirected and unweighted networks, i.e., the dolphin network and a network generated by the Barabási-Albert (BA) model. We used eight other undirected and unweighted networks, of which four are empirical networks and the other four are model networks, in our statistical analyses.
We downloaded the five empirical networks, including the dolphin network described in the main text, from the KONECT repository [61] at http://konect.cc. The other four empirical networks are as follows:

### v3

@@ II.6 Weighted averaging
We find that the additional optimization of node weights leads to substantially smaller approximation error, $\varepsilon$, than the unweighted averages examined in the previous sections. For comparison, across all dynamics and networks, $\varepsilon$ is, on average, $46$ times smaller when we optimize both the node set and the individual node weights than when we only optimize for the node set and constrain all $a_{i}=1$ (Supplementary Section S16; see Table S22).
Intriguingly, this advantage does not extend to the transferability across different dynamics. Specifically, although weight optimization improves performance on the training dynamics, it exhibits no meaningful reduction in $\varepsilon$ against the test dynamics ($p=0.591$; see Table S23). Our supposition is that while the composition of $S$ is determined mainly by the network structure, the specific weight to assign to each node in $S$ does depend on the dynamics, and therefore weight optimization against a specific training dynamics offers no advantage when applied to the test dynamics. Consequently, although optimizing node weights drastically improves the approximation of $\overline{x}$, it is only advantageous if we know the system’s dynamics, but cannot help in case of unknown dynamics.
@@ III Discussion
As our understanding of complex system dynamics advances [3, 1, 28, 29, 30, 31, 32], a crucial bottleneck that we continue to face is our limited empirical access to the states of their multitude of relevant parameters. This bottleneck hinders the development of new theoretical analyses and empirical validation of existing theories on dynamics. A similar challenge in biology and medicine has been that it is often difficult to track every single metabolite, protein, or cell. In that context, the identification of biomarkers [33, 34] has offered dramatic breakthroughs, allowing us to gain insight into the state of the entire system by just tracking a small set of indicative parameters. Our work here envisions an equivalent opportunity in the context of complex networks. It shows that, by tracking a mere $\lfloor\ln N\rfloor$ of sentinel nodes, we can approximate the average state of the network, making monitoring dynamics scalable even for large networks.
@@
We have found that the detected sentinels are largely agnostic to the model of the dynamics describing the system’s internal mechanisms. This result indicates that one can transfer the learned set of nodes from one dynamics to infer the state of the system under another dynamics. Such transfer learning is especially useful for systems with unknown dynamics, a rather common challenge in the context of complex networked systems [35, 36, 37, 38, 39]. To take advantage of transfer learning, one first needs to select a relevant model of the dynamics, numerically simulate the system under this model, and use our optimization method to extract the sentinel node set. This set, found under the selected model, will then help approximate the system’s state even if it is driven by some other, potentially unknown, dynamics.
Our method is heavily based on numerical analysis, and therefore its complexity depends on the network size $N$. In Supplementary Section S17, we analyze this dependence and show that both the dynamic simulation and the node set optimization scale as $O(N^{2})$. Therefore, even for large networks, the computational bottleneck is non-restrictive, and the entire analysis pipeline is scalable. In fact, our full analysis, provided in the Supplementary Information, covers networks with as many as $N\approx 10^{5}$ nodes, a clear testament to the method’s scalability. To further capitalize on this potential advantage, future work should aim to extract a master model, i.e., a nonlinear system of equations whose sentinel nodes optimally transfer to as many other types of dynamics as possible. We also seek to broaden the family of dynamics that our optimization can cover, for example, by considering dynamical systems with multiple equilibria or oscillations. Should the optimized sentinel node sets depend on the specific equilibrium point? Can oscillatory dynamics also be monitored in such a parsimonious fashion? The implicit assumption of monostable dynamics captures an important limitation of the present study.
While our methodology shows good performance when tested against numerically simulated dynamics, our long-term objective is to apply it to empirical multivariate time-series data. Our preliminary results using fMRI data offer an encouraging step in this direction. However, at this stage, the optimized sentinel node sets provide only a modest improvement over the random or degree-preserving alternatives. This limitation primarily stems from the highly constrained nature of the available data. Future applications employing richer datasets, specifically data with reliable information pertaining to the underlying causal network, are expected to enhance the method’s effectiveness.
@@
Our work differs from other approaches for approximating the population activity of or efficiently observing a network by attempting to accurately estimate $\overline{x}$. First, there are methods to select a small number of sentinel nodes for constructing early warning signals for anticipating regime shifts [16, 17, 18, 19, 20], but there is no a priori reason to consider that these sentinel nodes are also good at representing $\overline{x}$. Second, the theory of network observability allows us to determine the node set, observation of which enables one to reconstruct the system’s complete activity [40]. The method has also been applied for selecting sentinel nodes with which to construct early warning signals [18]. However, this network observability concerns inference of the initial condition, i.e., the activity of each node at time 0. In fact, $\overline{x}$ in the equilibrium or the time course of it, rather than its initial value, would be of practical interest in applications. Functional observers aim to approximate the equilibrium of any given weighted linear average of $x_{i}$, including $\overline{x}$ as a special case, using a linear dynamics of small order [21, 22]. Functional observability methods are mathematically founded and more versatile than our method in that they can approximate a wider variety of quantities. However, they are limited to linear and fixed dynamics. In contrast, our method accommodates nonlinear dynamics, tracks state transitions as a function of the system’s control parameters, and—most importantly—yields insight even when the underlying dynamical equations are unknown. Together, these features offer a substantial practical advantage.
Our approach also differs from earlier efforts to identify sentinel sites or individuals during the initial stages of an epidemic, particularly under realistic constraints such as limited observations of infected nodes and incomplete network information. While our algorithm assumes access to the network structure, it can tolerate partial knowledge (see Supplementary Section S13). Extending the method to transient dynamics—such as those observed in the early phases of an outbreak—is left for future work. One possible direction is to assume that the degree of each node is known, because this is often easier to obtain than the full network structure. A synthetic network can then be generated using the configuration model [2], allowing the algorithm to be applied in this approximate setting. This procedure is conceptually related to the degree-preserving random node sets explored in this paper and may offer a viable approximation of epidemic dynamics at early stages.
Existing reduction methods often place much weight on the hub nodes. Specifically, GBB [9] uses a degree weighted average, and DART [10, 11] weighs all nodes based on the principal eigenvector, both favoring observing $x_{i}^{*}$ of the hubs [41, 42]. Therefore, in degree-heterogeneous networks, the observed global state of the system is dominated by a small set of highly central nodes. This phenomenon resembles our formalism, which also achieves observability via a limited set of sentinels. However, these methods and our sentinel node approximation select very different nodes. While the nodes dominating GBB and DART tend to be the hubs, our sentinels tend to avoid the hubs. This discrepancy is rooted in the fact that the two approaches aim to observe distinct characteristics of the system. GBB and DART seek a theoretical closure for the dynamical equations, which is indeed driven by the behavior of the hubs. Our method, in contrast, seeks to observe the average $\overline{x}$, which is best represented by majority of small- and intermediate-degree nodes, rather than by the network outliers.
This contrast touches on the key trade-off between the GBB/DART approach and our machine learning-based observability. GBB and DART are derived from theoretical principles, and hence offer full interpretability, namely, we understand their selected observable $\overline{x}^{\prime}$. Our approach is based on optimization, thus sacrificing the interpretability for efficiency. Indeed, without running our full algorithm, one cannot construct the sentinel node set $S$, just from theoretical principles. This discussion captures the general balance between theory and machine learning-based approaches, trading accuracy and efficiency for opaqueness. We acknowledge that an important limitation of the present study is the lack of theoretical understanding of why our sampling strategy balancing different node degrees performs well. Despite these limitations, however, we did extract several interpretable insights relating to the expected features of sentinel nodes, such as the avoidance of hubs and the set-typical degree distribution. As commonly observed with machine learning frameworks, these insights provide a retrospective understanding of $S$, but cannot truly contribute to a priori predicting it. As the use of our proposed framework increases, we hope to obtain a richer mechanistic understanding of the sentinel nodes, which can help further reduce the search space for $S$. Analyzing weakly connected networks by theoretical means together with subtyping networks may help in achieving such goals.
@@
Our framework accommodates directed and weighted networks, noisy dynamics, and, the common scenario in which the system’s driving equations are unknown. Furthermore, without essential changes, our method allows any weighted sum of $\{x_{1}^{*},\ldots,x_{N}^{*}\}$ as the target quantity to be approximated. Together, these features make our method versatile enough for potential applications. For example, in ecosystem monitoring, our algorithm could help focus monitoring efforts on a small number of key species [43]. As we have shown, the selected sentinel species may not be the most abundant or well-connected species [44, 45]. Assuming that an estimated network of species interactions is available, which is often the case [46], one can derive a small number of sentinel species by simulating a ground truth using a mutualistic species dynamics model on the measured network. Then, the populations of the chosen sentinel species are expected to provide a good approximation of population dynamics of the entire ecosystem.
@@
From a methodological perspective, our drastic reduction from $N$ to just a small number of nodes is rooted in a combined toolbox that mixes theoretical modeling with machine learning techniques. The theoretical models, captured within the low-dimensional dynamical systems derived by GBB, DART, or other methods, represent our analytical insights into the system’s internal mechanisms. Such insights, however, are limited by the prohibitive complexity of the system. This is precisely where the machine learning component helps complement the theoretical approach, lending its computational power to help simplify the scale of the problem. This synergy between theoretical insights and the predictive power of optimization and learning methods represents, in our view, one of the main avenues towards predicting, observing, and influencing complex network behavior.
@@ IV.1 Node Selection
Given all $x_{i,\ell}^{*}$ values (with $i\in\{1,\ldots,N\}$ and $\ell\in\{1,\ldots,L\}$) computed, we determine $S$ by combinatorial simulated annealing. Specifically, we first initialize $S$ by selecting $n$ nodes uniformly at random from the network and calculate the approximation error for $S$, which we denote by $\varepsilon(S)$. We arbitrarily set $n=\lfloor\ln N\rfloor$, where $\lfloor\cdot\rfloor$ is the floor operation, unless we state otherwise. Then, in each $h$th iteration of the algorithm, we select an $i$th node uniformly at random from $S$ and replace it with a $j$th node that does not belong to $S$ and is chosen uniformly at random. The tentative new node set is given by $S^{\prime}=S\setminus\{i\}\cup\{j\}$. We compute $\varepsilon$ for $S^{\prime}$, which we denote by $\varepsilon(S^{\prime})$. We accept $S^{\prime}$ as the new $S$ according to the Metropolis criterion, i.e., with probability $1$ if $\varepsilon(S^{\prime})<\varepsilon(S)$ and with probability $q=\exp\{-[\varepsilon(S^{\prime})-\varepsilon(S)]/t\}$ otherwise. The normalization factor $t=t_{0}/\ln(h+e-1)$ decreases as $h\in\{1,\ldots,h_{\max}\}$ increases. Therefore, an $S^{\prime}$ with a larger approximation error than $S$ is less likely to be accepted as $h$ increases, encouraging convergence to a local minimum of $\varepsilon$. We set $t_{0}=10$. We also set $h_{\max}=50N$ and confirmed that this number of iterations was sufficient for reaching a local minimum of $\varepsilon$.
@@ IV.2 Dynamical system models
where $x_{i}$ represents the probability that the $i$th node is infectious, $\lambda$ is the infection rate, and $\mu$ is the recovery rate [54]. The second term on the right-hand side expresses the rate at which the $j$th node infects the $i$th node. We set $\mu=1$ without loss of generality; changing $(\lambda,\mu)$ to $(c\lambda,c\mu)$ with a constant $c$ is equivalent to scaling the time by a factor of $c$, so it does not affect the equilibrium. We use $\lambda$ as the control parameter in the range $[0,1]$. One obtains $x_{i}=0$ $\forall i$ when $\lambda$ is below a value called the epidemic threshold. Above the epidemic threshold, all $x_{i}^{*}$ (<1<1) values are positive given a positive initial value. The parameter $\lambda$ has the same role as $D$ in the other dynamics models.
@@
To compute equilibrium values of $x_{i}$ at each control parameter value, we initially set each $x_{i}$ to a model-dependent standard value (coupled double-well: $x_{i}=1$, SIS: $x_{i}=0.01$, mutualistic species: $x_{i}=0.001$, gene-regulatory: $x_{i}=2$). Then, we solved the ODEs using the implicit Adams method provided by the deSolve package for R [55]. We used a total simulation time of $T=15$, which we found was sufficient to allow each of the above dynamics on each network described below to relax to a point where no further change was noticeable. We used the value of each $x_{i}$ at $T=15$ as $x_{i,\ell}^{*}$.
@@ IV.3 Networks
We chose fifteen empirical and five model networks to test our method. All networks were coerced to be undirected, simple (i.e., no self- or multi-edges), and unweighted. We used the largest connected component.
@@
We used eighteen other networks as well as the dolphin and BA networks in the ANOVA analyses. See SI section S1 for the description of these eighteen networks.
@@ IV.4 Optimization of node weights
In section II.6, we approximate $\overline{x}_{\ell}$ with a weighted average of the node states in $S$, $\overline{x}^{\prime}_{\ell}=\sum_{i=1;i\in S}^{N}a_{i}x_{i,\ell}^{*}$. The GBB and DART assign weights to all the nodes in the network. In contrast, here we determine weights only for the nodes in $S$ by a quadratic programming optimization.
@@
where $\bm{1}=(1,\ldots,1)^{\top}$. We use the qpOASES algorithm [58] to solve for $\bm{a}$ via the ROI package [59] in R. In an iteration of our algorithm, we first tentatively update $S$ by carrying out one step of the algorithm for the unweighted averaging, optimize the node weights, $\bm{a}$, given the updated $S$ by solving Eq. (9), and calculate $\varepsilon$ for the updated $S$ and optimized $\bm{a}$. We then compare the obtained $\varepsilon$ with the value of $\varepsilon$ before updating $S$ and optimizing $\bm{a}$, and apply the Metropolis criterion to determine whether or not we adopt the tentatively updated $S$ and $\bm{a}$. We iterate this process as described in Sec. IV.1.
@@
Data availability
Empirical networks used in this study are available from https://networks.skewed.de or the original sources as described in the SI. All other data was produced by simulation and is available at https://zenodo.org/uploads/15848170 [60].
Code availability
Optimization, simulation, and analysis code used in this study is available at
https://github.com/ngmaclaren/observe [61].
Acknowledgments
B.B. was supported by the Israel Science Foundation (grant no. 499/19), the Israel-China ISF-NSFC joint research program (grant no. 3552/21), and by the VATAT grant for data science research. N. Masuda was supported by the Japan Science and Technology Agency (JST) Moonshot R&D (under grant no. JPMJMS2021), the National Science Foundation (under grant no. 2052720), and JSPS KAKENHI (under grant nos. JP 21H04595, 23H03414, and 24K14840). This work was performed in part at Center for Computational Research, the State University of New York at Buffalo.
@@ Acknowledgments
Competing interests
The authors declare no competing interests.
@@
Observing network dynamics through sentinel nodes
@@
In the main text, we have shown results from two undirected and unweighted networks, i.e., the dolphin network and a network generated by the Barabási-Albert (BA) model. We used eighteen other undirected and unweighted networks, of which fourteen are empirical networks and the other four are model networks, in our statistical analyses. For each network described in this section, we removed edge weights and the multiplicity of edges, coerced the network to be undirected if it was not so already, and selected the largest connected component.
We downloaded five empirical networks, including the dolphin network described in the main text, from the KONECT repository [62] at http://konect.cc. Besides the dolphin network, the other four empirical networks are as follows:
@@
We downloaded an additional ten networks from the Netzschleuder repository [67] as follows:
FlyBi
A network of protein interactions for the fruit fly Drosophila melanogaster [68]. A node is a protein. Two nodes are adjacent if the two proteins were observed to interact in a biochemical analysis. The network has $N=2,705$, $M=8,458$, $\overline{k}=6.25$, and CV of the degree equal to $1.76$.
Reactome
A network of human protein interactions [69]. Nodes represent proteins. Edges represent interactions. The network has $N=5,973$, $M=145,778$, $\overline{k}=48.81$, and CV of the degree equal to $1.39$.
Route views
A network of routine message passing between internet routers [70]. A node is a collection of internet protocol addresses belonging to the same entity. Two nodes are adjacent if border gateway protocol (BGP) messages passed between the two entities. This network has the largest number of nodes in the data set and was collected on January 2, 2000. The network has $N=6,474$, $M=12,572$, $\overline{k}=3.88$, and CV of the degree equal to $6.44$.
Spanish words
A network of word adjacencies drawn from the novel “Don Quixote” by Miguel de Cervantes [71]. Nodes are words. Two nodes are adjacent if one word appeared next to the other in the novel. The network has $N=11,558$, $M=43,050$, $\overline{k}=7.45$, and CV of the degree equal to $7.77$.
FOLDOC
A network of cross-references in an online encyclopedia, the Free On-line Dictionary of Computing [72]. A node is encyclopedia entry, which is a webpage. Two nodes are adjacent if there is a hyperlink from one page to the other. The network has $N=13,356$, $M=91,471$, $\overline{k}=13.70$, and CV of the degree equal to $1.13$.
Tree of life
A protein interaction network, similar to those used above [73]. We chose the network with the largest number of nodes available in the Netzschluder repository. The network has $N=16,415$, $M=440,122$, $\overline{k}=53.62$, and CV of the degree equal to $1.62$.
English words
A network of word associations [74]. A node is an English word. Two nodes are adjacent if a study participant gave one word as a response to another, stimulus word. The network has $N=23,132$, $M=297,094$, $\overline{k}=25.69$, and CV of the degree equal to $1.74$.
Enron
A network drawn from the Enron email corpus [75]. Nodes are email addresses. Nodes are adjacent if at least one email was sent from one address to the other. The network has $N=33,696$, $M=180,811$, $\overline{k}=10.73$, and CV of the degree equal to $3.50$.
Marker Cafe
An online social network [76]. A node is a user of the Marker Cafe social network. Two nodes are adjacent if one user is in the other’s “circle.” The network has $N=69,317$, $M=1,644,794$, $\overline{k}=47.46$, and CV of the degree equal to $3.72$.
Prosper
A network of loans in the online marketplace prosper.com [77]. A node is a user. Two nodes are adjacent if one user lent the other money. The network has $N=89,171$, $M=3,329,970$, $\overline{k}=74.69$, and CV of the degree equal to $1.87$.